# https://shengqi77.github.io/RLR-AT.github.io/ Long-range Turbulence Mitigation: A Large- scale Dataset and A Coarse-to-fine Framework

### v1

@@ 4.1 Frequency-aware Reference Frame Construction
Relationship between FRF and Temp Avg. We like to comment on the relationship between Temp Avg and proposed FRF, which is established through the parameter . The hyper-parameter in Eq. (2) determines the sensitivity of weight function to frequency. We can observe that when , the weight is one for all values of frequency, and the Eq. (3) can be simplified as:
@@ 4.2 Low-rank Tensor Distortion Refinement
Directly estimating multiple variables is difficult. To solve , we adopt the alternating minimization scheme [33] to solve the Eq. (8) with respect to each variable. Please refer to the supplementary material for the algorithm details.
@@ 5.1 Datasets and Experimental Settings
Comparison Methods. We compare CDSP with (1) conventional turbulence removal methods: TurbRecon [37], SG [35], CLEAR [4] and NDL [58]; (2) deep learning based methods: TurbNet [39], TSR-WGAN [29], PiRN [27], NDIR [32] and TMT [56]. For a fair and comprehensive comparison, considering that the experimental datasets consist of turbulence sequences in static scenes, we categorize the above methods into two groups: static-scene dependent and static-scene independent, and they are further classified based on input frames into multi-frame based and single-frame based. Though TMT and TurbRecon are static-scene independent methods, we employ their codes and pre-train models designed for static scenes to ensure a fair comparison. All methods incorporate deblurring effects, with NDIR and NDL using their default deblurring approach [45]. The codes and pre-train models of all methods are provided by the original authors.
@@ 6 Conclusion
Our work focuses on long-range turbulence mitigation. We construct a long-range turbulence dataset (RLR-AT). RLR-AT is the farthest and largest-scale turbulence dataset. We propose a coarse-to-fine framework for long-range turbulence mitigation, which cooperates the dynamic and the static priors. We propose a frequency-aware reference frame for better registration. We propose a low-rank tensor refinement model to refine the registration error with details preserving. Extensive experiments demonstrate the proposed method outperforms SOTA methods on different datasets.

### v2

@@ 4.1 Frequency-aware Reference Frame Construction
Relationship between FRF and Temp Avg. We further discuss the relationship between Temp Avg and FRF, which is established through the parameter . The in Eq. (2) decides the sensitivity of weight function to frequency. When , the weight is one for all intensities, and the Eq. (3) can be simplified as:
@@ 4.2 Low-rank Tensor Distortion Refinement
To solve , we adopt the alternating minimization scheme [33] to solve the Eq. (8) for each variable. Please refer to the appendix for the solution.
@@ 5.1 Datasets and Experimental Settings
Comparison Methods. We compare CDSP with (1) conventional turbulence removal methods: TurbRecon [37], SG [35], CLEAR [4] and NDL [58]; (2) deep learning based methods: TurbNet [39], TSR-WGAN [29], PiRN [27], NDIR [32] and TMT [56]. For a fair comparison, considering that the experimental datasets consist of static scenes, we categorize the methods into two groups: static-scene dependent and static-scene independent, and they are further classified based on input frames into multi-frame based and single-frame based. We employ codes and pre-train models of TMT and TurbRecon designed for static scenes to ensure a fair comparison. All methods incorporate deblurring effects, with NDIR and NDL using their default deblurring approach [45].
@@ 6 Conclusion
Our work focuses on long-range turbulence mitigation. We construct a long-range turbulence dataset (RLR-AT). We propose a coarse-to-fine framework for long-range turbulence mitigation, which cooperates the dynamic and the static priors. We propose a frequency-aware reference frame for better registration. We propose a low-rank tensor refinement model to refine the registration error with details preserving. Extensive experiments demonstrate the proposed method outperforms SOTA methods on different datasets.